40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 11930     Entry time: Wed Jan 13 18:36:00 2016
Author: gautam 
Type: Update 
Category: LSC 
Subject: restoration of green beat electronics 

In preparation for tonight's work, I did the following:

On the PSL table:

  • Powered the RF amplifiers for the green beat signal on
  • Reconnected the outputs of the Green beat PDs to the RF amplifiers
  • Restored wiring in the fiber box such that both IR beats go to the frequency counter.

At the IOO Rack area:

  • Restored wiring to the frequency counter module such that the IR beats from both arms go to the respective channels
  • Partially cleaned up the setup used for measuring AUX laser frequency noise - moved the SR785 to the X end along with one SR560 so that we can measure the end PDH OLTF
  • Brought the HP network analyzer back to the control room so that we can view the green beatnotes.

At the X-end:

  • Turned the function generator used for PDH locking back on
  • Checked that the AUX laser diode current is 1.90 A, and the crystal temperature is ~47.5 degrees, both of which I think are "good" values from our AUX laser frequency noise measurements
  • Did some minor manual alignment of the PZT mirrors

At the Y-end:

  • Restored the BNC connection from the PDH box to the laser's "FAST" control input. The long BNC cable used for the PLL is still running along the Y-arm, I will clean this up later.

Having done all this, I checked the green transmission levels for both arms (PSL green shutter closed, after running ASS to maximize IR transmission). GTRY is close to what I remember (~0.40) while the best I could get GTRX to is ~0.12 (I seem to remember it being almost double this value - maybe the alignment onto the beat PD has to be improved?). Also, the amplitudes of the beatnotes on the network analyzer are ~-50dBm, and I seem to remember it being more like -25dBm, so maybe the alignment on the PD is the issue? I will investigate further in the evening. It remains to measure the OLTF of the X-end PDH as well.

ELOG V3.1.3-